# DATransNet: Dynamic Attention Transformer Network for Infrared Small Target Detection

### v2

@@ II-A Dynamic Attention Transformer (DATrans)
Figure 2: $D$ is obtained by performing edge detection convolution on $T$ and unfolding the output.
@@
Figure 3: The various heads in DATrans employ the edge convolution to get $D_{\sum}$. Each head’s key and value are derived from $D_{\sum}$, while the query is generated from the flattened input.
@@
In the various heads of DATrans, we apply different edge detection convolutions from Fig. 2 to $T$, utilizing varying dilation ratios. The outputs from these convolutions are then reshaped into $D^{n}\in\mathbb{R}^{8c_{inp}\times wh}$. Additionally, the input $T$ is reshaped to form $O\in\mathbb{R}^{c_{inp}\times wh}$. The token $O$ serves as the query, and the local difference matrix $D^{n}\in\mathbb{R}^{8c_{inp}\times wh}$ functions as both the key and the value.
@@
where $W^{n}_{Q}\in\mathbb{R}^{\frac{c_{o}}{m}\times c_{inp}}$, $W^{n}_{K}\in\mathbb{R}^{\frac{8c_{inp}}{m}\times 8c_{inp}}$, $W^{n}_{V}\in\mathbb{R}^{\frac{8c_{inp}}{m}\times 8c_{inp}}$ are the reshaped weights of convolutions, $n$ denotes dilation ratios, and $m$ denotes the number of heads. We utilize the $Q^{n}\in\mathbb{R}^{\frac{c_{o}}{m}\times wh}$ and $K^{n}\in\mathbb{R}^{\frac{8c_{inp}}{m}\times wh}$ to produce the attention matrix $M^{n}\in\mathbb{R}^{\frac{c_{o}}{m}\times\frac{8c_{inp}}{m}}$. $V^{n}\in\mathbb{R}^{\frac{8c_{inp}}{m}\times wh}$ denotes the value of the transformer. The result $Out^{n}\in\mathbb{R}^{\frac{c_{o}}{m}\times wh}$ is defined as follows:

### v3

@@ II-A Dynamic Attention Transformer (DATrans)
Figure 2: $D^{n}$ is obtained by performing edge detection convolution with dilation ratio of n on $T$ and unfolding the output.
@@
Figure 3: The overall structure of DATrans.
@@
In the various heads of DATrans, we apply different edge detection convolutions from Fig. 2 to $T$, utilizing varying dilation ratios. The outputs from these convolutions are then reshaped into $D^{n}\in\mathbb{R}^{8c_{inp}\times wh}$ where $n$ is the dilation ratio. Additionally, the input $T$ is reshaped to form $O\in\mathbb{R}^{c_{inp}\times wh}$. The token $O$ serves as the query, and the local difference matrix $D^{n}\in\mathbb{R}^{8c_{inp}\times wh}$ functions as both the key and the value.
@@
where $W^{n}_{Q}\in\mathbb{R}^{\frac{c_{o}}{m}\times c_{inp}}$, $W^{n}_{K}\in\mathbb{R}^{\frac{8c_{inp}}{m}\times 8c_{inp}}$, $W^{n}_{V}\in\mathbb{R}^{\frac{8c_{inp}}{m}\times 8c_{inp}}$ are the reshaped weights of convolutions and $m$ denotes the number of heads. We utilize the $Q^{n}\in\mathbb{R}^{\frac{c_{o}}{m}\times wh}$ and $K^{n}\in\mathbb{R}^{\frac{8c_{inp}}{m}\times wh}$ to produce the attention matrix $M^{n}\in\mathbb{R}^{\frac{c_{o}}{m}\times\frac{8c_{inp}}{m}}$. $V^{n}\in\mathbb{R}^{\frac{8c_{inp}}{m}\times wh}$ denotes the value of the transformer. The result $Out^{n}\in\mathbb{R}^{\frac{c_{o}}{m}\times wh}$ is defined as follows: